OpenStudy (anonymous):

hi I was wondering if anyone could help me with finding the explicit rule for -2,-8,-18,-32,-50

OpenStudy (aum):

-2,-8,-18,-32,-50 Take the differences: -8+2, -18+8, -32+18, -50+32 = -6, -10, -14, -18 Take differences again: -10+6, -14+10, -18+14 = -4, -4, -4 (Differences are constant the second time). Therefore, the relationship is quadratic: a * n^2 where n = 1, 2, 3, 4 and 5. put n = 1, an^2 = a = -2 Therefore -2 * n^2 is the explicit rule where n = 1, 2, 3, 4 and 5.

OpenStudy (aum):

\(\large -2n^2\)

OpenStudy (anonymous):

Thank you so much really appreciate it :)